# Survey in Characterization of Semantic Change

### v4

@@ 5.6 Semantic Change Characterization
While the first equation only measured some notion of distance between the representations, for characterization we first apply the type function on the representation, to then compare. For example in Buechel2016FeelingsFT we have a orientation characterization, e.g.,
@@ 5.7 Framework Evaluation
To illustrate the framework’s practical aspects, we investigate two corpora: SEMCOR (Miller1993ASC) and MASC [^4] https://anc.org/data/masc/corpus/. SEMCOR is a sense-tagged corpus of  200,000 words created in 1993. MASC is a manually annotated sub-corpus of  500,000 words of modern American English released around 2013.
@@
SEMCOR and MASC are corpora manually annotated using WordNet senses. We obtain annotations for the poles of change from two lexical resources: ChainNet, a WordNet extension linking senses by metaphor and metonymy, and SentiWordNet, which provides a positive/negative score for every sense in WordNet. This step allows us to skip sense inference and focus on demonstrating the framework. For automatic WordNet annotation, we suggest Navigli2017WordSD. For work on annotating data for the poles of change independently of a lexical resource, we suggest Sa2024SemanticCC. From SEMCOR and MASC, we computed occurrences of ‘heart’ and ‘plane,’ two well-known words that have changed semantically.
@@ 6.1 Dimension
A limitation for approaches based on dictionaries is the temporal delay between a neologism’s first appearance and its inclusion in dictionaries. Another limitation is the presence of metaphorical senses at the same level as literal ones, increasing the complexity of distinguishing between ’dimension’ and ’relation.’ Some authors prefer to group these two poles into one (Tang2013SemanticCC). For topic- and graph-based approaches, the limitation is that interpreting the meaning of a topic or a graph branch is not always clear and can require expert intervention.
@@ 6.2 Relation
Outside of this LSC survey, a vast body of work proposes methods to predict if a word in a sentence has a metaphorical sense (Heintz2013AutomaticEO; Shutova2013UnsupervisedMI; Hovy2013IdentifyingMW). These approaches analyze texts but don’t consider temporal information, so they cannot determine if a word has undergone metaphorization over time. Some works claim to detect the metaphorization of words (e.g., (Tredici2016TracingMI)), but they rely on the general hypothesis that any sense change implies a metaphorical change, which still needs to be demonstrated.
For metonymization, we could not find studies that performed metonymy identification in a comparative setting. However, related recent studies investigate metonymy usage with various methods for material relation identification (Nastase2012LocalAG; BaldiniSoares2019MatchingTB; AlexMathews2021ImpactOT). We expect that future investigations can extend these approaches to detect word metonymization.
@@
Tang2013SemanticCC mainly rely on dictionaries of heuristics to infer metaphoricity, which served for qualitative analysis, but still require more extensive validation. A single work used ground-truth annotations for metaphors (Schlechtweg2017GermanIF), but did not reported a final precision of the proposed methodology.
@@ 6.3 Orientation
Interest in orientation analysis is increasing with the wide adoption of social networks. Orientation is usually determined using a feeling function ($\mathfrak{f}$) that relies on seed words and the assumption that these seed words don’t vary over time. As we illustrated, word meanings evolve and can differ between communities. Most works found in this survey addressed the change in orientation for particular words Xiao2023HaveTC; MillerLewis2021WordsDF.
The example of ‘sick‘ shows that a word can have two opposite orientations depending on the context. Creating corpora that include all these variations with statistical significance is a complex task. The annotation process can suffer from many biases, both from annotators and data. Selected seed words may not generalize across all contexts or long time spans, and a series of misleading results can occur, as presented in Antoniak2021BadSE.
@@
While there is a huge body of work on sentence-level sentiment analysis, fewer studies infer sentiment at the word level (Guo2021AnOO), and only a small fraction of those do so in a comparative manner, as we have assessed in this survey. While works in topics provide lexical entries that could be used similar to what was done in embeddings, this was not investigated so far.
An important outcome of this survey is that there is some harmonization in how change in feeling is defined and applied. However, we did not observe a harmonized approach to error analysis for the feeling function. Some works adopt a binary orientation function (positive, negative), while others use multidimensional orientations (valence, arousal, dominance) (Buechel2016FeelingsFT). Further study is needed to analyze which function ($\mathfrak{f}$) is most suitable for characterizing change.
A converging point from the many approaches analyzed is the difficulty of evaluating and comparing results because, for most tasks, no gold standard exists. Many papers have raised this problem (Schlechtweg2017GermanIF), but some initiatives to create reference datasets are underway. Another point is multilingualism; the vast majority of experiments use only English documents, making it difficult to generalize methods to other languages. The recent emergence of new-generation language models has opened up new possibilities for identifying and characterizing semantic changes, but this area of research is still under-explored.

### v5

@@ 5.6 Semantic Change Characterization
While the first equation only measured some notion of distance between the representations, for characterization we first apply the type function on the representation, to then compare. For example in Buechel et al. (2016) we have a orientation characterization, e.g.,
@@ 5.7 Framework Evaluation
To illustrate the framework’s practical aspects, we investigate two corpora: SEMCOR (Miller et al., 1993) and MASC [^4] https://anc.org/data/masc/corpus/. SEMCOR is a sense-tagged corpus of  200,000 words created in 1993. MASC is a manually annotated sub-corpus of  500,000 words of modern American English released around 2013.
@@
SEMCOR and MASC are corpora manually annotated using WordNet senses. We obtain annotations for the poles of change from two lexical resources: ChainNet, a WordNet extension linking senses by metaphor and metonymy, and SentiWordNet, which provides a positive/negative score for every sense in WordNet. This step allows us to skip sense inference and focus on demonstrating the framework. For automatic WordNet annotation, we suggest Navigli et al. (2017). For work on annotating data for the poles of change independently of a lexical resource, we suggest de Sá et al. (2024). From SEMCOR and MASC, we computed occurrences of ‘heart’ and ‘plane,’ two well-known words that have changed semantically.
@@ 6.1 Dimension
A limitation for approaches based on dictionaries is the temporal delay between a neologism’s first appearance and its inclusion in dictionaries. Another limitation is the presence of metaphorical senses at the same level as literal ones, increasing the complexity of distinguishing between ’dimension’ and ’relation.’ Some authors prefer to group these two poles into one (Tang et al., 2013). For topic- and graph-based approaches, the limitation is that interpreting the meaning of a topic or a graph branch is not always clear and can require expert intervention.
@@ 6.2 Relation
Outside of this LSC survey, a vast body of work proposes methods to predict if a word in a sentence has a metaphorical sense (Heintz et al., 2013; Shutova and Sun, 2013; Hovy et al., 2013). These approaches analyze texts but don’t consider temporal information, so they cannot determine if a word has undergone metaphorization over time. Some works claim to detect the metaphorization of words (e.g., (Tredici et al., 2016)), but they rely on the general hypothesis that any sense change implies a metaphorical change, which still needs to be demonstrated.
For metonymization, we could not find studies that performed metonymy identification in a comparative setting. However, related recent studies investigate metonymy usage with various methods for material relation identification (Nastase et al., 2012; Soares et al., 2019; Mathews and Strube, 2021). We expect that future investigations can extend these approaches to detect word metonymization.
@@
Tang et al. (2013) mainly rely on dictionaries of heuristics to infer metaphoricity, which served for qualitative analysis, but still require more extensive validation. A single work used ground-truth annotations for metaphors (Schlechtweg et al., 2017), but did not reported a final precision of the proposed methodology.
@@ 6.3 Orientation
Interest in orientation analysis is increasing with the wide adoption of social networks. Orientation is usually determined using a feeling function ($\mathfrak{f}$) that relies on seed words and the assumption that these seed words don’t vary over time. As we illustrated, word meanings evolve and can differ between communities. Most works found in this survey addressed the change in orientation for particular words Xiao et al. (2023); Miller-Lewis et al. (2021).
The example of ‘sick‘ shows that a word can have two opposite orientations depending on the context. Creating corpora that include all these variations with statistical significance is a complex task. The annotation process can suffer from many biases, both from annotators and data. Selected seed words may not generalize across all contexts or long time spans, and a series of misleading results can occur, as presented in Antoniak and Mimno (2021).
@@
While there is a huge body of work on sentence-level sentiment analysis, fewer studies infer sentiment at the word level (Guo et al., 2021), and only a small fraction of those do so in a comparative manner, as we have assessed in this survey. While works in topics provide lexical entries that could be used similar to what was done in embeddings, this was not investigated so far.
An important outcome of this survey is that there is some harmonization in how change in feeling is defined and applied. However, we did not observe a harmonized approach to error analysis for the feeling function. Some works adopt a binary orientation function (positive, negative), while others use multidimensional orientations (valence, arousal, dominance) (Buechel et al., 2016). Further study is needed to analyze which function ($\mathfrak{f}$) is most suitable for characterizing change.
A converging point from the many approaches analyzed is the difficulty of evaluating and comparing results because, for most tasks, no gold standard exists. Many papers have raised this problem (Schlechtweg et al., 2017), but some initiatives to create reference datasets are underway. Another point is multilingualism; the vast majority of experiments use only English documents, making it difficult to generalize methods to other languages. The recent emergence of new-generation language models has opened up new possibilities for identifying and characterizing semantic changes, but this area of research is still under-explored.
@@ Acknowledgments
Use of AI tools
The production of this manuscript used AI tools solely to improve grammar and readability, as English is not the authors’ native language.